I have no idea how to compute this. My txtbook doesn't help. please help me!!!
Compute
P(E | F)
and P(F | E).
(to three decimal places.)
P(E) = 0.70, P(F) = 0.30, P(E ∩ F) = 0.24
P(E | F)=
P(F | E)=

Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

I have no idea how to compute this. My txtbook doesn't help. please help me!!!
Compute
P(E | F)
and P(F | E).
(to three decimal places.)
-----
Your text probably gives you a "Conditional Probability Rule"
which says P(A and B) = P(A|B)*P(B)
-------------------

P(E) = 0.70, P(F) = 0.30, P(E ∩ F) = 0.24
P(E | F)= [P(E and F)]/P(F) = 0.24/0.30 = 4/5
--------------------
P(F | E)= [P(F and E)/P(E) = 0.24/0.70 = 0.3429
==============
Cheers,
Stan H.
==============

Answer by ewatrrr(24785)

 
Hi,
Note: P(A|B) = P(A ∩ B)/P(B)
P(E) = 0.70, P(F) = 0.30, P(E ∩ F) = 0.24
P(E | F)= P(E ∩ F)/P(F) = .24/.30
P(F | E)= P(F ∩ E)/P(E) = .24/.70